Evaluate the integral of e^(tanx) sec² x dx


1
Expert's answer
2021-03-30T08:11:17-0400

Since "(tan x)'=sec^2 x", use a substitution: "t = tan x", "dt = sec^2 x dx"


"\int e^{tan x} sec^2 x dx = \int e^t dt =e^t +C"

Substitute "t = tan x" back into result:


"e^{tan x} +C"